OpenStudy (anonymous):

What is the solution of 5 - 2x > -3 or -3x + 9 < -6?

OpenStudy (anonymous):

@Olive99 Can you help me solve this one?

OpenStudy (anonymous):

solve for each one

OpenStudy (anonymous):

see < or > as = :)

OpenStudy (anonymous):

First, for 5-2x>-3 subtract 5 both side -2x>-8 and you need to divide both side by -2 P.S. :Because you are dividing both sides by a negative number, do not forget to change or flip the inequality sign. Tht's the rule. I want you to try solving it from here. For the second one, I would subtract 9 both side and do the same procedures as the above one. Again, when dividing by a negative number, flip the sign. Try it and if you are lost, just hit me up and I will solve it for you. I want you to try it yourself so you can do better on he tests.

OpenStudy (anonymous):

5 - 2x > -3 or -3x + 9 < -6 5 - 2x = -3 or -3x +9 =-6 ( does this look easier? )

OpenStudy (anonymous):

5 - 2x = -3 or -3x +9 =-6 -2x= -8 or -3x = -15 x = 4 or x= 5 is your results

OpenStudy (anonymous):

but don't stop at that

OpenStudy (anonymous):

Please do not use the equation sign unless you think you will remember to change it to inequality sign and to flip the sign when dividing by the negative number

OpenStudy (anonymous):

nah

OpenStudy (anonymous):

but it easier to look :)

OpenStudy (anonymous):

but remember to change it back to inequality signs

OpenStudy (anonymous):

as you divided with negative numbers :v let's me finish my words women :v

OpenStudy (anonymous):

@Olive99 How did you get -2x>-8 ?

OpenStudy (anonymous):

you NEED TO CHANGE YOUR (>) OR (<)

OpenStudy (anonymous):

5-2x>-3 add 5 both side -2x>-8 divide by -2 both side x<4 -3x+9<-6 subtract 9both side -3x<-15 divide by -3 both side x>5

OpenStudy (anonymous):

oh well, she got it. It just when you divide the negative number ( the -x) you have to change > into < or opposite

OpenStudy (anonymous):

unless is +x then you don't have to worry about the changing :)

OpenStudy (anonymous):

@Olive99 Thank you for explaining it more! I'm so new to this.

OpenStudy (anonymous):

So the solution is x>5?

OpenStudy (anonymous):

nah

OpenStudy (anonymous):

you have 2 answers since you have 2 different equations :) so 1 of them would be x>5 and the other would be x<4 :)

OpenStudy (anonymous):

Oh okay! Thank you!!

OpenStudy (anonymous):

ur welcome, and give the nice girl Oliver over there a medal please :3 cuz she's nice :3